OpenStudy (anonymous):

The diameter of a circular pool is 18 feet. What is the approximate area of the pool? A. 56.52 square feet B. 113.04 square feet C. 254.34 square feet D. 1017.36 square feet

OpenStudy (anonymous):

C

OpenStudy (ranga):

Area = \(\pi * r^2\); Radius \(r = \frac{diameter}{2}\)

OpenStudy (anonymous):

yup that^ :3

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

np he deserves the medal :p give it to him xD